OpenStudy (anonymous):

what is the equation of the line perpendicular to 9x + 2y = 18 and passes through the point (1, 4)?

OpenStudy (anonymous):

eq of line perp. to 9x + 2y = 18 is 2x-9y=k but our reqd line passes through (1,4) hence 2-36=k or k=-34 hence the line is 2x-9y =-34